How do I solve?
A light ray moving at a 54.3 deg angle in water (n 1.33) hits a boundary with oil (n 1.52). At what angle is it refracted in the oil? (Water n 1.33, Air n 1.00) (Unit - deg)
Answer #1

Using Snell's law

n1*sin(theta1) = n2*sin(theta2)

1.33*sin(54.3) = 1.52*sin(theta2)

sin(theta2) = 1.33*0.81208353/1.52 = 0.71057

Theta2 = 45.25 degree in oil